# AI Feedback Enhances Community-Based Content Moderation through Engagement with Counterarguments

### v1

@@ Abstract
Today, social media platforms are significant sources of news and political communication, but their role in spreading misinformation has raised significant concerns. In response, these platforms have implemented various content moderation strategies. One such method, Community Notes on X, relies on crowdsourced fact-checking and has gained traction, though it faces challenges such as partisan bias and delays in verification. This study explores an AI-assisted hybrid moderation framework in which participants receive AI-generated feedback —supportive, neutral, or argumentative —on their notes and are asked to revise them accordingly. The results show that incorporating feedback improves the quality of notes, with the most substantial gains resulting from argumentative feedback. This underscores the value of diverse perspectives and direct engagement in human-AI collective intelligence. The research contributes to ongoing discussions about AI’s role in political content moderation, highlighting the potential of generative AI and the importance of informed design.
@@ Introduction
The emergence of social media platforms has transformed how users consume and share information. Studies show that these platforms have become primary news sources for many individuals [1, 2] and are widely used by politicians as campaign tools [3]. This shift presents challenges for platform owners in monitoring and verifying the information being shared. The widespread dissemination of misleading content and misinformation has become a primary concern for social media platforms. This issue is especially pronounced during significant socio-political events, such as the 2016 U.S. presidential election [4] and the Brexit referendum [5]. The spread of false information during times of crisis can seriously mislead the public. For example, the rapid circulation of misinformation about COVID-19 led the WHO’s Director-General to describe the situation as an ”infodemic” [6]. Another striking case is the January 6th, 2021, attack on the U.S. Capitol, where social media posts helped spread the misinformation that incited violence [7]. In response to these concerns, social media platforms have implemented various content moderation strategies, including both community-based and automated approaches.
@@
Community Notes rely on the wisdom of the crowd to identify misleading content and leverage the diversity of users on the platform to produce moderated outcomes. The process involves displaying specific posts to users in the Community Notes program and asking them to write a note that provides context for those posts. Users are also asked to indicate whether they believe the post is misleading and to provide reasoning for their judgment. Subsequently, other users rate the helpfulness of these notes. Initially, notes with high helpfulness ratings appeared beneath the posts in users’ timelines. However, early analyses revealed strong partisan behaviour. Yasseri and Menczer (2023) found that the network connecting note writers and raters mirrors the highly polarised friendship network on X [12]. Similarly, Allen et al. (2022) showed that users are more likely to write notes on posts by counter-partisans, label them as misleading, and rate notes from counter-partisans as unhelpful, while rating notes from co-partisans as helpful [13].
@@
However, recent advances in generative AI offer promising solutions to the limitations of existing automated content moderation strategies [21, 22]. Research shows that Large Language Models (LLMs) can effectively persuade individuals on political issues [23] and generate fact-checking messages that are more persuasive than those produced by human experts [24]. In addition, several studies demonstrate that LLMs can accurately assess the credibility of news sources [25] and verify factual claims [26, 27, 28], with their accuracy improving when fine-tuned on augmented data [29]. In a recent example, De et al. (2025) fine-tuned an LLM on the Community Notes dataset to generate what they call “supernotes,” which were rated significantly more helpful than notes written by Community Notes users for the same posts [30]. These findings suggest that LLMs are not only capable of understanding complex content but may also outperform human-written moderation in certain cases. However, studies have also found that purely automated approaches can be less effective in the presence of social influence [31]. Moreover, people tend to place considerable trust in LLMs, even when the models make mistakes, raising concerns about overreliance on AI-generated content [32].
@@
An increasingly supported view in the literature is that the most effective content moderation strategies combine human oversight with AI assistance [33]. This hybrid approach leverages the complementary strengths of humans and machines, enabling moderation that is both more nuanced and scalable [34]. Schmitt et al. (2024) suggest that models integrating human judgment with AI can outperform either component alone [35]. However, even within hybrid systems, users often overestimate the accuracy of LLMs, placing trust in AI-generated assessments even when they are incorrect—a phenomenon known as automation bias or machine heuristics [36, 37]. Moreover, some Human-AI frameworks risk reinforcing users’ existing beliefs, potentially deepening echo chambers and contributing to political polarisation [38]. To mitigate these risks, such systems must be carefully designed to avoid unintentionally amplifying bias and division on social media platforms.
@@
Despite these risks, political diversity itself can be a valuable asset in improving community-based moderation efforts. One of the key strengths of Community Notes is its ability to harness ideological diversity—even polarisation—to produce more balanced and neutral notes [39]. A similar dynamic has been observed on Wikipedia, where research attributes the platform’s success to the variety of viewpoints among editors, along with its structured policies and guidelines [40, 41, 42, 43]. Shi et al. (2019), for instance, found that higher levels of polarisation among editors correlate with higher-quality articles, especially on politically sensitive topics [42]. Moreover, studies suggest that exposure to cross-partisan messages can reduce both affective and issue-based polarisation [44, 45]. These findings indicate that, when properly structured, political diversity not only does not hinder collaboration but also can even enhance it. In a recent study, researchers paired two participants to collaboratively write Community Notes. They found that duos produced more helpful notes than individuals, and that politically diverse groups were especially effective when evaluating posts with Republican-leaning content [46]. These results further support the idea that political diversity can strengthen the note-writing process in Community Notes, provided it is applied in a way that is both effective and efficient.
@@
Building on these insights, this study proposes a human-AI collaboration framework that uses AI-generated feedback to simulate political diversity in content moderation. The aim is to combine the efficiency of AI with the interpretive strengths of human contributors, while ensuring that humans retain final decision-making authority. We investigate whether AI-generated feedback can emulate cross-partisan collaboration and enhance the quality of contributions on the Community Notes platform. Our focus is on how users revise their notes in response to different types of AI-generated feedback. Specifically, this study addresses the following research questions:
@@
To investigate these research questions, we conducted an experiment modelled after the Community Notes process. Participants () were shown a post and asked to write a note providing additional context. Each note was then submitted to a GPT-4 language model, which was prompted to generate one of three types of feedback: argumentative, supportive, or neutral.
Argumentative feedback offers counterpoints to the original note, supportive feedback reinforces the note’s arguments, and neutral feedback simply restates its content. While all feedback was generated by the AI model, we randomly told some participants that the feedback was written by a human expert. This manipulation allows us to examine the roles of human favouritism [47] and automation bias [36] in shaping participants’ responses.
Participants were then asked to revise their notes in light of the feedback. In a second experiment (), we crowdsourced evaluations of note quality before and after revision. Self-identified Republicans and Democrats rated the notes on a 0–10 scale based on their helpfulness. Figure 1 provides a schematic overview of the experimental workflow.
@@ Results
To measure engagement, we introduce the Feedback Acceptance rate (), a metric that quantifies the extent to which participants incorporated the provided feedback into their final submissions. Specifically, is defined as the semantic similarity score between the feedback and the revised note (see Methods for details). To evaluate whether engagement with feedback leads to higher-quality revisions, we use as a predictor in ordinal logistic regression models. The dependent variables capture whether the revised note has declined, was unchanged, or improved according to the Democrat raters () and Republican raters (). The results (Table 1) show that is a strong and consistent predictor of note improvement as perceived by both partisan groups. A comparison of odds ratios further indicates that has a slightly stronger effect on improvement ratings from Democrats than from Republicans.
| odds ratio | odds ratio | |||
| 3.581(0.604) | <0.001 | 2.498(599) | <0.001 | |
| Observations | 893 | 893 | ||
| Pseudo | 0.019 | 0.009 | ||
| Deviance | 1872.86 | 1924.63 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -936.43 | -962.32 | ||
| Akaike Inf. Crit. | 1878.86 | 1930.63 | ||
@@
To better understand the factors that influence this engagement metric, we analyse the relationship between and three key variables: the type of post (supporting either Democratic or Republican ideology), the participant’s self-identified partisanship, and the type of feedback received. The results from the Ordinary Least Squares (OLS) regression are shown in Table 2.
@@
The analysis shows that users who received argumentative feedback were less likely to incorporate it into their final note than those who received supportive feedback. Additionally, Democratic participants were more likely than Republicans to use the feedback in their final notes. However, the type of post did not have a significant effect on this behaviour.
@@
However, analysing the type of post in isolation may not provide the complete picture. We expect individuals who share the political ideology of the post’s author to respond differently when moderating content compared to those with opposing views. To investigate this, we introduce a binary alignment variable that captures the relationship between a participant’s self-identified partisanship and the post’s stance: co-partisan if the participant and the post share the same ideology, and cross-partisan if they do not. A post is classified as co-partisan when it supports Democrats (Republicans) and the participant identifies as a Democrat (Republican); otherwise, it is classified as cross-partisan.
| Ordinary Least Squares (OLS) Regression | Dependent variable: | |
| Coefficient | ||
| Feedback (Argue) | -0.094(0.008) | <0.001 |
| Feedback (Support) | -0.052 (0.008) | <0.001 |
| Post (D) | 0.001 (0.007) | 0.880 |
| Participant (D) | 0.021 (0.007) | 0.001 |
| Constant | 0.850 (0.007) | <0.001 |
| Observations | 893 | |
| 0.143 | ||
| Adjusted | 0.139 | |
| Residual Std. Error | 0.097 (df = 888) | |
| F Statistics | 36.899 (df = 4; 888) | |
@@
To examine how political alignment affects engagement with feedback, we conducted an OLS regression with as the dependent variable, including an interaction term between co-partisanship and feedback type. The results of this analysis are shown in Table 3. The findings indicate that when participants share the same political affiliation as the post, they are less likely to incorporate argumentative feedback. A similar pattern emerges for supportive feedback: participants are also less likely to adopt supportive feedback language when the post aligns with their political views.
| Ordinary Least Squares (OLS) Regression | Dependent variable | |
| Coefficient | ||
| Co-partisan × Neutral | -0.019 (0.011) | 0.090 |
| Cross-partisan × Argue | -0.102 (0.012) | <0.001 |
| Co-partisan × Argue | -0.104 (0.011) | <0.001 |
| Cross-partisan × Support | -0.059 (0.011) | <0.001 |
| Co-partisan × Support | -0.064 (0.011) | <0.001 |
| Constant | 0.871 (0.008) | <0.001 |
| Observations | 893 | |
| 0.135 | ||
| Adjusted | 0.130 | |
| Residual Std. Error | 0.098 (df = 887) | |
| F Statistics | 27.711 (df = 5; 887) | |
@@
The second treatment condition manipulated the labelled source of the feedback. Specifically, we investigate whether participants respond differently depending on whether they believe the feedback was generated by a human expert or an AI agent. To explore this, we replicate the previous regression analysis on two separate subsets of the data: one in which feedback was attributed to an AI agent, and another in which it was attributed to a human expert. Figure 3 presents the marginal predicted values of across the different treatment conditions for both feedback source labels. While the overall patterns are consistent across conditions, we observe a slight increase in when the feedback is labelled as AI-generated, specifically for cross-partisan posts. This pattern holds for both argumentative and supportive feedback. These results suggest a marginal preference for AI-labelled feedback when participants engage with content that opposes their political views. However, the differences are not statistically significant (e.g., predicted for AI in cross-partisan × argumentative = 0.777 vs. 0.762 for human expert, ; predicted for AI in cross-partisan × supportive = 0.822 vs. 0.803 for human expert, ), indicating that the labelled source of feedback does not meaningfully influence user behaviour in this context.
@@
To construct the full model and evaluate the effect of feedback type conditional on participant engagement, we estimated an ordinal logistic regression including both the main effects of feedback type and its interaction with . The dependent variables are and , which reflect the improvement of the notes according to Democrat and Republican raters, respectively. This model allows us to assess whether argumentative or supportive feedback is more likely to lead to improvements in perceived helpfulness when participants actively engage with the feedback, and whether these effects differ across political subgroups.
@@
The results of this analysis are presented in Table 4. The findings indicate that once participants incorporate feedback into their notes, both argumentative and supportive feedback increase the likelihood of producing a more helpful note, as rated by both Democratic and Republican evaluators. However, argumentative feedback shows a larger effect size than supportive feedback, suggesting that when participants engage with feedback that challenges their views and includes counterarguments, the resulting notes are more likely to be of higher quality.
| odds ratio | odds ratio | |||
| Feedback (Argue) | -1.210 (1.362) | 0.374 | -1.944 (1.346) | 0.148 |
| Feedback (Support) | -0.583 (1.228) | 0.635 | -0.765 (1.238) | 0.536 |
| Feedback (Neutral) × | 2.643 (0.987) | 0.007 | 1.534 (1.001) | 0.125 |
| Feedback (Argue) × | 3.713 (1.370) | 0.006 | 3.817 (1.331) | 0.004 |
| Feedback (Support) × | 3.662 (1.086) | <0.001 | 2.237 (1.079) | 0.038 |
| Observations | 893 | 893 | ||
| Pseudo | 0.027 | 0.011 | ||
| Deviance | 1857.37 | 1921.71 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -928.68 | -960.86 | ||
| Akaike Inf. Crit. | 1871.37 | 1935.71 | ||

### v3

@@ Abstract
Today, social media platforms are significant sources of news and political communication, but their role in spreading misinformation has raised significant concerns. In response, these platforms have implemented various content moderation strategies. One such method, Community Notes (formerly Birdwatch) on X (formerly Twitter), relies on crowdsourced fact-checking and has gained traction. However, it faces challenges such as partisan bias and delays in verification. This study explores an AI-assisted hybrid moderation framework in which participants receive AI-generated feedback, supportive, neutral, or argumentative, on their notes and are asked to revise them accordingly. The results show that incorporating feedback improves the quality of notes, with the most substantial gains resulting from argumentative feedback. This underscores the value of diverse perspectives and direct engagement in human-AI collective intelligence. The research contributes to ongoing discussions about AI’s role in political content moderation, highlighting the potential of generative AI and the importance of informed design.
@@ Introduction
The emergence of social media has transformed how users consume and share information. Studies show that these platforms have become primary news sources for many individuals [1, 2] and are widely used by politicians as campaign tools [3]. This shift presents challenges for platform owners in monitoring and verifying the information being shared. The widespread dissemination of misleading content and misinformation has become a primary concern for social media. This issue is especially pronounced during significant socio-political events, such as the 2016 U.S. presidential election [4] and the Brexit referendum [5]. The spread of false information during times of crisis can seriously mislead the public. For example, the rapid circulation of misinformation about COVID-19 led the WHO’s Director-General to describe the situation as an ”infodemic” [6]. Another striking case is the January 6th, 2021, attack on the U.S. Capitol, where social media posts helped spread the misinformation that incited violence [7]. In response to these concerns, social media platforms have implemented various content moderation strategies, including both community-based and automated approaches.
@@
In contrast, Community Notes rely on the wisdom of the crowd to identify misleading content and leverage the diversity of users on the platform to produce moderated outcomes. Users write a note on a misleading post to provide additional context. Other users can then evaluate these notes by rating their helpfulness. In the pilot version, notes with high helpfulness ratings appeared beneath the posts in users’ timelines. However, early analyses revealed strong partisan behaviour. Yasseri and Menczer (2023) found that the network connecting note writers and raters mirrors the highly polarised friendship network on X [12]. Similarly, Allen et al. (2022) showed that users are more likely to write notes on posts by counter-partisans, label them as misleading, and rate notes from counter-partisans as unhelpful, while rating notes from co-partisans as helpful [13].
@@
However, recent advances in generative AI offer promising solutions to the limitations of existing automated content moderation strategies [21, 22]. Research shows that Large Language Models (LLMs) can effectively persuade individuals on political issues [23] and generate fact-checking messages that are more persuasive than those produced by human experts [24]. In addition, several studies demonstrate that LLMs can accurately assess the credibility of news sources [25] and verify factual claims [26, 27, 28], with their accuracy improving when fine-tuned on augmented data [29]. In a recent example, De et al. (2025) fine-tuned an LLM on the Community Notes dataset to generate what they call “supernotes”, which were rated significantly more helpful than notes written by Community Notes users for the same posts [30]. Similarly, recent work has shown that AI can extract fact-checking information from user comments on a post to produce contextual notes that are rated more helpful than Community Notes [31]. These findings suggest that LLMs are not only capable of understanding complex content but may also outperform human-written moderation in certain cases. However, studies have also found that purely automated approaches can be less effective in the presence of social influence [32]. Moreover, people tend to place considerable trust in LLMs, even when the models make mistakes, raising concerns about overreliance on AI-generated content [33].
@@
An increasingly supported view in the literature is that the most effective content moderation strategies combine human oversight with AI assistance [34]. This hybrid approach leverages the complementary strengths of humans and machines, enabling moderation that is both more nuanced and scalable [35]. AI systems can process vast volumes of content at speeds unattainable for human moderators. Humans, in turn, provide contextual understanding, moral reasoning, and sensitivity to cultural and political nuances that remain challenging for AI to replicate. Schmitt et al. (2024) suggest that models integrating human judgment with AI can outperform either component alone [36]. However, even within hybrid systems, users often overestimate the accuracy of LLMs, placing trust in AI-generated assessments even when they are incorrect—a phenomenon known as automation bias or machine heuristics [37, 38]. Moreover, some Human-AI frameworks risk reinforcing users’ existing beliefs, potentially deepening echo chambers and contributing to political polarisation [39]. To mitigate these risks, such systems must be carefully designed to avoid unintentionally amplifying bias and division on social media platforms.
One proposed strategy is the deliberate introduction of “biased AI”. Typically, AI assistants are trained to generate outputs that users rate highly [23, 40]. However, in hybrid systems where both humans and AI are intended to contribute their respective strengths, overly compliant AI may lead users to over-rely on the system and engage less of their own cognitive resources. This phenomenon, referred to as metacognitive laziness, has been observed across different tasks [41, 42]. Prior research shows that strategic biases can reactivate human critical thinking and improve decision-making [43]. Introducing bias in controlled ways may enhance human productivity within Human-AI systems. For instance, a recent study on collaboration between humans and biased AI in evaluating the veracity of news headlines found that participants performed better when exposed to opposing views. However, they regarded the partisan AI as less trustworthy [44].
@@
Furthermore, introducing AI with perspectives that differ from those of human collaborators can inject political diversity into the moderation process. Political diversity itself can be a valuable asset in improving community-based moderation efforts. One of the key strengths of Community Notes is its ability to harness ideological diversity—even polarisation—to produce more balanced and neutral notes [45]. A similar dynamic has been observed on Wikipedia, where research attributes the platform’s success to the variety of viewpoints among editors, along with its structured policies and guidelines [46, 47, 48, 49]. Shi et al. (2019), for instance, found that higher levels of polarisation among editors correlate with higher-quality articles, especially on politically sensitive topics [48]. Moreover, studies suggest that exposure to cross-partisan messages can reduce both affective and issue-based polarisation [50, 51]. These findings suggest that, when properly structured, political diversity not only does not hinder collaboration but can also enhance it. In a recent study, researchers paired two participants to write Community Notes collaboratively. They found that duos produced more helpful notes than individuals, and that politically diverse groups were especially effective when evaluating posts with Republican-leaning content [52]. These results further support the idea that political diversity can strengthen the note-writing process in Community Notes, provided it is applied in a way that is both effective and efficient.
While diversity can strengthen the collective intelligence of teams, its benefits depend on how effectively members can collaborate. Collective intelligence emerges not only from the diversity of perspectives but also from the team’s capacity to integrate resources, combine individual knowledge into joint problem-solving, and implement solutions in practice [53, 54]. Research shows that one of the strongest predictors of collective intelligence is the quality of collaboration processes among team members [55]. These processes include effective interaction mechanisms such as coordination and cooperation, which are essential for enabling groups to reach outcomes that exceed individual contributions [56]. Importantly, this principle applies not only to human teams but also to hybrid human–AI collaborations [22].
@@
Building on these insights, this study proposes a human-AI collaboration framework that uses AI-generated feedback to simulate political diversity in content moderation. The aim is to combine the efficiency of AI with the interpretive strengths of human contributors, while ensuring that humans retain final decision-making authority. We investigate whether AI-generated feedback can emulate cross-partisan collaboration and enhance the quality of contributions on the Community Notes platform. Our focus is on how users revise their notes in response to different types of AI-generated feedback. This work contributes to the growing body of research on the dynamics of Human–AI collaboration, especially in sensitive domains such as content moderation. Specifically, this study addresses the following research questions:
@@
To investigate these research questions, we conducted an experiment modelled after the Community Notes process. Participants () were shown a post and asked to write a note providing additional context. Each note was then submitted to the GPT-4 language model, which was prompted to generate one of three types of feedback: argumentative, supportive, or neutral.
Argumentative feedback offers counterpoints to the original note, supportive feedback reinforces the note’s arguments, and neutral feedback simply restates its content. We intentionally did not prompt the LLM to adopt a partisan stance (i.e., Democrat or Republican); instead, the prompt focused solely on the arguments in the note, thereby avoiding stereotyping of either the AI or the participants.
While the language model generated all the feedback, we randomly instructed about half of the participants in each group that a human expert had written the feedback. This manipulation allows us to examine the roles of human favouritism [57] and automation bias [37] in shaping participants’ responses.
Participants were then asked to revise their notes in light of the feedback they received. In a second experiment (), we crowdsourced evaluations of note quality before and after revision. Self-identified Republicans and Democrats rated the notes on a 0–10 scale based on their helpfulness. Figure 1 provides a schematic overview of the experimental workflow.
The argumentative AI is designed to function as a form of “biased AI”, providing counterarguments to the participants’ initial notes. This design aimed to encourage participants to leverage their full cognitive resources and to consider opposing viewpoints. Consistent with prior research [44], participants were generally less likely to engage with argumentative AI feedback. However, among participants who did engage with the AI feedback, argumentative feedback yielded the highest chance of improvement.
Overall, our results highlight the potential of AI to assist humans in sensitive tasks such as content moderation. They further illuminate the mechanisms of collective hybrid intelligence, demonstrating that the productivity of Human-AI teams depends critically on the extent to which humans actively engage with the AI and the kind of feedback AI provides. Additionally, our findings suggest that Human-AI collaboration can effectively simulate political diversity in the note-writing process, thereby enhancing the overall performance of Community Notes.
@@ Results
To measure engagement, we introduce the Feedback Acceptance rate (), a metric that quantifies the extent to which participants incorporated the provided feedback into their final submissions. Specifically, is defined as the semantic similarity score between the feedback and the revised note (see Methods for details). To evaluate whether engagement with feedback leads to higher-quality revisions, we use as a predictor in ordinal logistic regression models. The dependent variables capture whether the revised note has declined, was unchanged, or improved according to the Democrat raters () and Republican raters (). The results (Figure 3) show that is a strong and consistent predictor of note improvement as perceived by both partisan groups. A comparison of odds ratios further indicates that has a slightly stronger effect on improvement ratings from Democrats than from Republicans. The result of the regression is available in Table S1.
@@
To better understand the factors that influence this engagement metric, we analyse the relationship between and three key variables: the type of post (supporting either Democratic or Republican ideology), the participant’s self-identified partisanship, and the type of feedback received. The results from the Ordinary Least Squares (OLS) regression are shown in Figure 4 (a).
@@
The analysis shows that users who received argumentative feedback were less likely to incorporate it into their final note than those who received supportive feedback. Additionally, Democratic participants were more likely than Republicans to use the feedback in their final notes. However, the type of post did not have a significant effect on this behaviour (see Table S2).
@@
However, analysing the type of post in isolation may not provide the complete picture. We expect individuals who share the political ideology of the post’s author to respond differently when moderating content compared to those with opposing views. To investigate this, we introduce a binary alignment variable that captures the relationship between a participant’s self-identified partisanship and the post’s stance: co-partisan if the participant and the post share the same ideology, and cross-partisan if they do not. A post is classified as co-partisan when it supports Democrats (or Republicans) and the participant identifies as a Democrat (or Republican); otherwise, it is classified as cross-partisan.
@@
To examine how political alignment affects engagement with feedback, we conducted an OLS regression using as the dependent variable, including an interaction term between co-partisanship and feedback type. The results, shown in Figure 4 (b), reveal no significant differences in how users incorporate feedback into their notes, regardless of whether they share the same political affiliation as the post. This holds for both supportive and argumentative feedback (See Table S3).
@@
The second treatment condition manipulated the labelled source of the feedback. Specifically, we investigate whether participants respond differently depending on whether they believe the feedback was generated by a human expert or an AI agent. To explore this, we replicate the previous regression analysis on two separate subsets of the data: one in which feedback was attributed to an AI agent, and another in which it was attributed to a human expert. Figure 5 presents the values of across the different treatment conditions for both feedback source labels. While the overall patterns are consistent across conditions, we observe a slight increase in when the feedback is labelled as AI-generated, specifically for cross-partisan posts. This pattern holds for both argumentative and supportive feedback. These results suggest a marginal preference for AI-labelled feedback when participants engage with content that opposes their political views. However, the differences are not statistically significant (e.g., for AI in cross-partisan × argumentative = 0.777 vs. 0.762 for human expert, ; for AI in cross-partisan × supportive = 0.822 vs. 0.803 for human expert, ), indicating that the labelled source of feedback does not meaningfully influence user behaviour in this context.
@@
To construct the complete model and evaluate the effect of feedback type, conditional on participant engagement, we estimated an ordinal logistic regression that included both the main effects of feedback type and its interaction with . The dependent variables are and , which reflect the improvement of the notes according to Democrat and Republican raters, respectively. This model enables us to determine whether argumentative or supportive feedback is more likely to lead to improvements in perceived helpfulness when participants actively engage with the feedback, and whether these effects vary across different political subgroups.
@@
The results of this analysis are presented in Figure 6. The findings indicate that once participants incorporate feedback into their notes, both argumentative and supportive feedback increase the likelihood of producing a more helpful note, as rated by both Democratic and Republican evaluators. However, argumentative feedback shows a larger effect size than supportive feedback, suggesting that when participants engage with feedback that challenges their views and includes counterarguments, the resulting notes are more likely to be of higher quality.